# Unraveling the Key Components of OOD Generalization via Diversification

### v2

@@ G.1 Experimental details
  • Step 2 (Constructing training data where $h^{\star}$, $h_{\mathrm{MLP}}$ and $h_{\mathrm{RN}}$ completely correlates): As presented in Tab. 10, in $D_{t}$ row, we select the data points as training set $D_{t}$ such that $h^{\star}$, $h_{\mathrm{MLP}}$ and $h_{\mathrm{RN}}$ agree. As shown in (Atanov et al., 2022) by adversarial splits, when two hypotheses correlate with each other (i.e., their labels are the same on training data), a neural network tends to converge to the hypothesis with higher AS. Thus, combining with the conditions in step 1 (i.e., Eq. 3), training MLP on $D_{t}$ with ERM shoudl converge to $h_{\mathrm{MLP}}$ and training ResNet on $D_{t}$ with ERM to $h_{\mathrm{RN}}$, which is illustrated in Tab. 2-Right.
@@
$h^{\star}\perp h_{\mathrm{ViT}}$ ViT 76.4±${}_{\pm 0.6}$
ResNet18 58.1±${}_{\pm 0.4}$
$h^{\star}\perp h_{\mathrm{RN}}$ ViT 68.1±${}_{\pm 0.3}$
ResNet18 79.0±${}_{\pm 0.5}$

### v3

@@ G.1 Experimental details
  • Step 2 (Constructing training data where $h^{\star}$, $h_{\mathrm{MLP}}$ and $h_{\mathrm{RN}}$ completely correlates): As presented in Tab. 10, in $D_{t}$ row, we select the data points as training set $D_{t}$ such that $h^{\star}$, $h_{\mathrm{MLP}}$ and $h_{\mathrm{RN}}$ agree. As shown in (Atanov et al., 2022) by adversarial splits, when two hypotheses correlate with each other (i.e., their labels are the same on training data), a neural network tends to converge to the hypothesis with higher AS. Thus, combining with the conditions in step 1 (i.e., Eq. 3), training MLP on $D_{t}$ with ERM should converge to $h_{\mathrm{MLP}}$ and training ResNet on $D_{t}$ with ERM to $h_{\mathrm{RN}}$, which is illustrated in Tab. 2-Right.
@@
$h^{\star}\perp h_{\mathrm{ViT}}$ ViT 76.4±0.6
ResNet18 58.1±0.4
$h^{\star}\perp h_{\mathrm{RN}}$ ViT 68.1±0.3
ResNet18 79.0±0.5